# Low-probability Tokens Sustain Exploration in Reinforcement Learning with Verifiable Reward

### v1

@@ 5.1 Experimental Setup
For evaluation, we assess model performance across five diverse mathematical reasoning benchmarks: AIME24 (MAA, ), AIME25 (MAA, ), MATH-500 (Hendrycks et al., 2021), OlympiadBench (He et al., 2024a), and Minerva Math (Lewkowycz et al., 2022). Following (Cui et al., 2025), we employ distinct decoding strategies based on the benchmark. For AIME24 and AIME25, which have smaller test sets, we use sampled decoding with a temperature of $0.6$ and generate $16$ independent responses per problem to obtain a robust performance estimate. For the remaining benchmarks, including MATH-500, OlympiadBench, and Minerva, we utilize greedy decoding to evaluate performance.
@@ 5.2 Results
As shown in Figure 3 and Table 1, Lp-Reg achieves state-of-the-art performance across five challenging mathematical reasoning benchmarks on both 14B and 32B model scales. On the Qwen3-14B model, on-policy Lp-Reg sets a new benchmark with an average accuracy of $60.17\%$, surpassing the next best method, 80/20, by $2.66\%$. Notably, Lp-Reg’s advantage is more pronounced on the newer Qwen3-14B base model compared to the older Qwen2.5-32B. We hypothesize that as base models improve, their capacity for nuanced, low-probability reasoning increases, creating a richer substrate for the emergence of valuable reasoning sparks, which Lp-Reg can then effectively protect and leverage.
Our experiments consistently show the superiority of on-policy training over off-policy methods across 14B and 32B scales. This is due to the inherent stability of on-policy updates, which avoid distribution shifts caused by mismatched data-sampling and training policies. Off-policy methods, such as Clip-Higher, often rely on importance sampling clipping, leading to instability. While competitive on Qwen2.5-32B, Clip-Higher’s performance drops on Qwen3-14B, highlighting its fragility. In contrast, Lp-Reg’s self-contained, policy-intrinsic regularization ensures its effectiveness in both on-policy and off-policy settings, unlike competing methods that are heavily reliant on off-policy importance sampling.
@@
Qwen2.5-32B-Base
GRPO (Shao et al., 2024a) (off.) 30.63 22.29 88.00 41.18 54.37 47.29
GSPO (Zheng et al., 2025) (off.) 33.33 22.29 87.60 48.53 55.56 49.46
Clip-Higher (Yu et al., 2025) (off.) 38.33 29.79 87.60 45.22 56.44 51.48
KL-Cov (Cui et al., 2025) (off.) 35.62 27.50 87.40 44.49 55.11 50.02
80/20 (Wang et al., 2025) (off.) 38.12 28.75 87.00 45.22 58.37 51.49
@@
Qwen3-14B-Base
GRPO (Shao et al., 2024a) (off.) 34.38 27.08 89.20 49.26 55.70 51.13
GSPO (Zheng et al., 2025) (off.) 41.46 34.58 88.60 50.74 59.85 55.05
Clip-Higher (Yu et al., 2025) (off.) 41.67 32.71 95.00 47.43 64.00 56.16
KL-Cov (Cui et al., 2025) (off.) 49.17 34.79 93.00 47.43 62.07 57.29
80/20 (Wang et al., 2025) (off.) 43.96 34.58 91.80 48.16 60.89 55.88
@@
Table 1: Main results on five mathematical reasoning benchmarks across two model scales. Our method, Lp-Reg, achieves the best average performance on both models. On-policy (on.) and off-policy (off.) training methods are highlighted with distinct colors.
Figure 3: Training dynamics on the Qwen3-14B-Base model. To best illustrate the performance differences, we compare the top-performing methods. Lp-Reg demonstrates superior and more stable performance throughout training. Full training dynamics for the Qwen2.5-32B model are available in Appendix A.1.
@@ 5.3 Ablation Study
Figure 4: Ablation studies for core components of Lp-Reg on the Qwen3-14B-Base model. The results confirm that targeting our noise filtering threshold $\tau$ is critical for stable performance. The adaptiveness of the min-p threshold is also shown to be beneficial over a fixed one.
We conduct a series of ablation studies to analyze the core components of Lp-Reg and validate our key design choices.
@@
Lp-Reg only protects tokens deemed meaningful by the proxy distribution ($\pi_{\text{proxy}}>0$). To test this, we remove the filter and fork all tokens below the noise threshold $\tau$ from contributing to gradient updating (Lp-Reg w/o $\tau$). Figure 4 shows that this leads to a catastrophic performance collapse and entropy explosion. This confirms that filtering is critical to ignore the extreme tail of the distribution, which consists of irrelevant noise that destabilizes training if regularized.
Dynamic vs. Fixed Threshold.
We conduct a comparison between the dynamic min-p noise threshold (Lp-Reg w/ dynamic $\tau$) and the fixed noise threshold (Lp-Reg w/ fixed $\tau$) in Section 4.1. As shown in Figure 4, the fixed threshold underperforms compared to the dynamic threshold, which we adopt as the default. However, it still significantly surpasses the standard GRPO. This indicates that while the core filtering principle is effective, the dynamic nature of min-p provides a more robust estimate of the model’s confidence across different contexts, better preserving genuine reasoning sparks.
We conduct further ablation studies on the high-entropy token regularization and reverse KL regularization. For detailed results and analysis, please refer to Appendix A.2.
@@ 6.1 Probability-Entropy Distribution of Exploratory Tokens
Figure 5: The word cloud statistics.
@@ 6.2 Sampling Dynamics of Exploratory Tokens
Figure 6: Probability–Entropy scatter plots of explorative tokens, displaying a random sample of $5\%$ of all data points.
@@ 6.3 Probabilistic Distinction between Reasoning Sparks and Noise
The results reveal a clear and consistent statistical distinction: across all training stages, the average next-token probability of meaningful exploratory tokens is persistently higher than that of irrelevant tokens. It can be attributed to the intrinsic confidence of LLMs (Nguyen et al., 2025; Xu et al., 2025; Fu et al., 2025). This persistent probabilistic gap provides the foundational justification for our Lp-Reg design. It suggests that while a perfect separation is not possible, a probability threshold $\tau$, as defined for our proxy distribution in Section 4.1, can serve as a principled filtering mechanism. By setting such a threshold, we can effectively filter out a substantial portion of the lowest-probability irrelevant tokens, which constitute destabilizing noise, while simultaneously retaining a majority of the valuable exploratory tokens that give rise to reasoning sparks. This allows Lp-Reg to focus its regularization on tokens that are more likely to be meaningful, providing a targeted and robust approach to preserving high-quality exploration.
@@ 7 Conclusion
In this work, we investigated the exploration collapse in Reinforcement Learning with Verifiable Rewards. We identified a key mechanism driving this failure: the systematic elimination of a class of valuable, low-probability exploratory tokens we term reasoning sparks. To address this, we introduced Low-probability Regularization (Lp-Reg), a method designed to selectively preserve these crucial exploratory pathways. Lp-Reg is founded on the key insight that within the low-probability range, meaningful exploratory tokens consistently exhibit higher average probabilities than semantically irrelevant tokens, whose low-probability appearances act as noise. By leveraging this statistical distinction to filter out irrelevant noise and regularizing the policy towards the remainder, our method effectively protects valuable low-probability tokens from being extinguished. This focus on exploration quality over quantity enables stable on-policy training for around 1,000 steps, resulting in a $2.66\%$ test accuracy improvement over baselines and underscoring the importance of preserving the reasoning sparks within the policy’s low-probability tail.
@@ Appendix A Details of Experiments
A.1 Further Training Dynamics
The training dynamics of Lp-Reg and other RLVR methods on the Qwen2.5-32B base model are presented in Figure 9. The results show that Lp-Reg maintains a comparable performance in test accuracy throughout the training process, underscoring the benefits of our low-probability token regularization strategy for preventing exploration collapse.
A.2 Further Ablation Study
Figure 10: Ablation study comparing low-probability token regularization versus high-entropy token regularization for Lp-Reg (on-policy) on the Qwen3-14B-Base model. The test datasets used are AIME-24 and AIME-25, with the average score reported over 16 test runs.
Low Probability vs. High Entropy.
Forward KL vs. Reverse KL.
Figure 11: Ablation study comparing the forward and reverse KL formulations for the Lp-Reg penalty on the Qwen3-14B-Base model. The results demonstrate the superiority of the forward KL, which uses the heuristic proxy distribution as a soft guide, over the reverse KL, which forces a strict imitation.
@@ A.2 Further Ablation Study
A.3 Hyperparameter Sensitivity Analysis
@@ B.3 Training Dynamics of Regularized Token
To better understand how Lp-Reg operates during training, we analyze the dynamics of the probability threshold $\delta_{\rho}^{\mathcal{B}}$ and the proportion of low-probability tokens subjected to regularization $\frac{|\pi_{\bm{\theta}}(o|\cdot)<\delta_{\rho}^{\mathcal{B}}\land\pi_{\text{proxy}}(o|\cdot)>0|}{|\pi_{\bm{\theta}}(o|\cdot)<\delta_{\rho}^{\mathcal{B}}|}$. As shown in Figure 13, the threshold $\delta_{\rho}^{\mathcal{B}}$ gradually decreases with training steps. In parallel, the regularization ratio also declines steadily. This trend suggests that as training progresses, the extreme low-probability range becomes increasingly dominated by irrelevant tokens, constituting what we term irrelevant noise. In parallel, the semantically meaningful exploratory tokens are lifted into higher-probability regions, thus requiring less regularization.
@@ B.4 Case Study
To further illustrate the effect of the filter applied on low-probability tokens, Figure 17 presents a case study of a model-generated response, where low-probability tokens are highlighted according to whether they were preserved or filtered. Tokens with probability greater than $\tau$ are those retained by the filter, while tokens with probability smaller than $\tau$ are suppressed. The preserved tokens include meaningful exploratory markers such as ”Then”, ”Wait”, which guide the reasoning trajectory, whereas the discarded set largely consists of relatively irrelevant tokens such as ”We”, ”also”, ”that”. This qualitative evidence complements our quantitative analysis, demonstrating that Lp-Reg effectively leverages the min-p transformation to distinguish between valuable low-probability exploratory tokens (reasoning sparks) and tokens that constitute destabilizing, irrelevant noise.
@@
Figure 17: An Example generated by Qwen3-14B-Base model trained by Lp-Reg from math reasoning.

### v2

@@ 5.1.3 Evaluation
For evaluation, we assess model performance across five diverse mathematical reasoning benchmarks: AIME24 (aime), AIME25 (aime), MATH-500 (hendrycks2021measuring), OlympiadBench (he-etal-2024-olympiadbench), and Minerva Math (lewkowycz2022solving). Following (cui2025entropymechanismreinforcementlearning), we employ distinct decoding strategies based on the benchmark. For AIME24 and AIME25, which have smaller test sets, we use sampled decoding with a temperature of $0.6$ and generate $16$ independent responses per problem to obtain a robust performance estimate. For the remaining benchmarks, including MATH-500, OlympiadBench, and Minerva, we utilize greedy decoding to evaluate performance.
@@ 5.2 Results
Figure 3: Continuous scaling over $3,000$ training steps, totaling $81,204$ GPU-hours, for Lp-Reg (on-policy) on the Qwen2.5-32B-Base model.
As shown in Figure 3, Lp-Reg enables a continuous reinforcement learning scaling for $3,000$ training steps, totaling $81,204$ GPU-hours on the Qwen2.5-32B-Base model. Furthermore, Figure 4 and Table 1 reveal that Lp-Reg achieves state-of-the-art performance across five challenging mathematical reasoning benchmarks on both 14B and 32B model scales. On the Qwen3-14B model, on-policy Lp-Reg sets a new benchmark with an average accuracy of $60.17\%$, surpassing the next best method, 80/20, by $2.66\%$. Notably, Lp-Reg’s advantage is more pronounced on the newer Qwen3-14B base model compared to the older Qwen2.5-32B in the first $1,000$ training steps. We hypothesize that as base models improve, their capacity for nuanced, low-probability reasoning increases, creating a richer substrate for the emergence of valuable reasoning sparks, which Lp-Reg can then effectively protect and leverage.
Our experiments consistently show the superiority of on-policy training over off-policy methods across 14B and 32B scales in Figure 4. This is due to the inherent stability of on-policy updates, which avoid distribution shifts caused by mismatched data-sampling and training policies. Off-policy methods, such as Clip-Higher, often rely on importance sampling clipping, leading to instability. While competitive on Qwen2.5-32B, Clip-Higher’s performance drops on Qwen3-14B, highlighting its fragility. In contrast, Lp-Reg’s self-contained, policy-intrinsic regularization ensures its effectiveness in both on-policy and off-policy settings, unlike competing methods that are heavily reliant on off-policy importance sampling.
@@
Qwen2.5-32B-Base ($800$ training steps)
GRPO (shao2024deepseekmathpushinglimitsmathematical) (off.) 30.63 22.29 88.00 41.18 54.37 47.29
GSPO (zheng2025groupsequencepolicyoptimization) (off.) 33.33 22.29 87.60 48.53 55.56 49.46
Clip-Higher (yu2025dapoopensourcellmreinforcement) (off.) 38.33 29.79 87.60 45.22 56.44 51.48
KL-Cov (cui2025entropymechanismreinforcementlearning) (off.) 35.62 27.50 87.40 44.49 55.11 50.02
80/20 (wang20258020rulehighentropyminority) (off.) 38.12 28.75 87.00 45.22 58.37 51.49
@@
Qwen3-14B-Base ($1,000$ training steps)
GRPO (shao2024deepseekmathpushinglimitsmathematical) (off.) 34.38 27.08 89.20 49.26 55.70 51.13
GSPO (zheng2025groupsequencepolicyoptimization) (off.) 41.46 34.58 88.60 50.74 59.85 55.05
Clip-Higher (yu2025dapoopensourcellmreinforcement) (off.) 41.67 32.71 95.00 47.43 64.00 56.16
KL-Cov (cui2025entropymechanismreinforcementlearning) (off.) 49.17 34.79 93.00 47.43 62.07 57.29
80/20 (wang20258020rulehighentropyminority) (off.) 43.96 34.58 91.80 48.16 60.89 55.88
@@
Table 1: Main results on five mathematical reasoning benchmarks across two model scales. On-policy (on.) and off-policy (off.) training methods are highlighted with distinct colors. We report the best score among the total training for each method.
Figure 4: Training dynamics on the Qwen3-14B-Base model. On-policy training exhibits better training stability and testing performance compared to off-policy training.
Figure 5: Training dynamics on the Qwen3-14B-Base model. To best illustrate the performance differences, we compare the top-performing methods. Lp-Reg demonstrates more stable performance throughout training.
@@ 5.3 Ablation Study
We conduct a series of ablation studies to analyze the core components of Lp-Reg and validate our key design choices for regularization in Equation 6, including low-probability regularization, noise filtering, negative sample regularization, and forward KL.
Importance of Low-Probability Token
Figure 7: Ablation study comparing low-probability token regularization versus high-entropy token regularization for Lp-Reg (on-policy) on the Qwen3-14B-Base model.
@@
Lp-Reg only protects tokens deemed meaningful by the proxy distribution ($\pi_{\text{proxy}}>0$). To test this, we remove the filter and fork all tokens below the noise threshold $\tau$ from contributing to gradient updating (Lp-Reg w/o $\tau$). Figure 8 shows that this leads to a catastrophic performance collapse and entropy explosion. This confirms that filtering is critical to ignore the extreme tail of the distribution, which consists of irrelevant noise that destabilizes training if regularized. We further conduct a comparison between the dynamic min-p noise threshold (Lp-Reg w/ dynamic $\tau$) and the fixed noise threshold (Lp-Reg w/ fixed $\tau$) in Section 4.1. As shown in Figure 8, the fixed threshold underperforms compared to the dynamic threshold, which we adopt as the default. However, it still significantly surpasses the standard GRPO. This indicates that while the core filtering principle is effective, the dynamic nature of min-p provides a more robust estimate of the model’s confidence across different contexts, better preserving genuine reasoning sparks.
Figure 8: Ablation studies for noise filtering of Lp-Reg (on-policy) on the Qwen3-14B-Base model. The results confirm that targeting our noise filtering threshold $\tau$ is critical for stable performance. The adaptiveness of the min-p threshold is also shown to be beneficial over a fixed one.
Effect of Negative Samples.
To dissect the contribution of our regularization strategy, we analyze the effects of applying the Lp-Reg penalty to negative samples versus both positive and negative samples. As illustrated in Figure 9, applying regularization to low-probability tokens in all samples (Regularize Pos. & Neg.) sustains exploration and yields continuous performance growth, outperforming standard GRPO (Regularize None). This finding validates that protecting low-probability tokens is essential for successful exploration in RLVR. Furthermore, the comparison reveals that concentrating the regularization solely on negative samples leads to a faster learning rate than applying it to both sample types. We attribute this to the primary function of Lp-Reg, which is to protect valuable low-probability tokens from being excessively penalized during training. These crucial tokens are far more prevalent in negative samples. Once these tokens are protected, extending the penalty to positive samples provides little additional exploratory advantage. Instead, it appears to introduce minor noise into the gradient updates, which, while not destabilizing, slows the overall training progress. Thus, we conclude that targeting negative samples exclusively is the most efficient and effective application of our method.
Figure 9: Ablation study comparing positive sample regularization and negative sample regularization for the Lp-Reg penalty (on-policy) on the Qwen3-14B-Base model. Negative sample regularization exhibits better performance.
Forward KL vs. Reverse KL.
Figure 10: Ablation study comparing the forward and reverse KL formulations for the Lp-Reg penalty (on-policy) on the Qwen3-14B-Base model. The results demonstrate the superiority of the forward KL, which uses the heuristic proxy distribution as a soft guide, over the reverse KL, which forces a strict imitation.
@@ 6.1 Probability-Entropy Distribution of Exploratory Tokens
Figure 11: The word cloud statistics from training samples at GRPO training step 1 on Qwen3-14B-Base.
@@ 6.2 Sampling Dynamics of Exploratory Tokens
Figure 12: Probability–Entropy scatter plots of five explorative tokens from training samples at training step 256 on Qwen3-14B-Base, displaying a random sample of $5\%$ of all data points. Detailed plots for individual tokens are analyzed in Appendix B.2
@@ 6.3 Probabilistic Distinction between Reasoning Sparks and Noise
The results reveal a clear and consistent statistical distinction: across all training stages, the average next-token probability of meaningful exploratory tokens is persistently higher than that of irrelevant tokens. It can be attributed to the intrinsic confidence of LLMs (nguyen2025turningheatminpsampling; xu2025adaptiveterminationmultiroundparallel; fu2025deepthinkconfidence). This persistent probabilistic gap provides the foundational justification for our Lp-Reg design. It suggests that while a perfect separation is not possible, a probability threshold $\tau$, as defined for our proxy distribution in Section 4.1, can serve as a principled filtering mechanism. By setting such a threshold, we can effectively filter out a substantial portion of the lowest-probability irrelevant tokens, which constitute destabilizing noise, while simultaneously retaining a majority of the valuable exploratory tokens that give rise to reasoning sparks. This allows Lp-Reg to focus its regularization on tokens that are more likely to be meaningful, providing a targeted and robust approach to preserving high-quality exploration.
@@ 7 Conclusion
In this work, we investigated the exploration collapse in Reinforcement Learning with Verifiable Rewards. We identified a key mechanism driving this failure: the systematic elimination of a class of valuable, low-probability exploratory tokens we term reasoning sparks. To address this, we introduced Low-probability Regularization (Lp-Reg), a method designed to selectively preserve these crucial exploratory pathways. Lp-Reg is founded on the key insight that within the low-probability range, meaningful exploratory tokens consistently exhibit higher average probabilities than semantically irrelevant tokens, whose low-probability appearances act as noise. By leveraging this statistical distinction to filter out irrelevant noise and regularizing the policy towards the remainder, our method effectively protects valuable low-probability tokens from being extinguished. This focus on exploration quality over quantity enables continuous on-policy scaling for around $3,000$ steps, resulting in an absolute $2.66\%$ test accuracy improvement over baselines and underscoring the importance of preserving the reasoning sparks within the policy’s low-probability tail.
@@ Appendix A Details of Experiments
A.1 Hyperparameter Sensitivity Analysis
@@ B.3 Training Dynamics of Regularized Token
To better understand how Lp-Reg operates during training, we analyze the dynamics of the probability threshold $\delta_{\rho}^{\mathcal{B}}$ and the proportion of low-probability tokens subjected to regularization $\frac{|\pi_{\bm{\theta}}(o|\cdot)<\delta_{\rho}^{\mathcal{B}}\land\pi_{\text{proxy}}(o|\cdot)>0|}{|\pi_{\bm{\theta}}(o|\cdot)<\delta_{\rho}^{\mathcal{B}}|}$. As shown in Figure 16, the threshold $\delta_{\rho}^{\mathcal{B}}$ gradually decreases with training steps. At the same time, the regularization ratio also declines steadily. This trend suggests that as training progresses, the extreme low-probability range becomes increasingly dominated by irrelevant tokens, constituting what we term irrelevant noise. It also explains the relatively worse performance using a fixed threshold $\tau=0.02$ than the dynamic min-p threshold $\tau$ in Equation 5, since a fixed threshold failed to regularize the lowest $1\%$ tokens while the dynamic one succeeded. Furthermore, the semantically meaningful exploratory tokens are lifted into higher-probability regions, thus requiring less regularization.
@@ B.4 Case Study
To further illustrate the effect of the filter applied on low-probability tokens, Figure 20 to Figure 22 present a case study of a model-generated response, where low-probability tokens are highlighted according to whether they were preserved or filtered. Tokens with probability greater than $\tau$ are those retained by the filter, while tokens with probability smaller than $\tau$ are suppressed. The preserved tokens include meaningful exploratory markers such as ”Then”, ”Wait”, which guide the reasoning trajectory, whereas the discarded set largely consists of relatively irrelevant tokens such as ”We”, ”also”, ”that”. This qualitative evidence complements our quantitative analysis, demonstrating that Lp-Reg effectively leverages the min-p transformation to distinguish between valuable low-probability exploratory tokens (reasoning sparks) and tokens that constitute destabilizing, irrelevant noise.
@@
Figure 20: An Example generated by Qwen3-14B-Base model trained by Lp-Reg from math reasoning. (Part 1)
Figure 21: An Example generated by Qwen3-14B-Base model trained by Lp-Reg from math reasoning. (Part 2)
Figure 22: An Example generated by Qwen3-14B-Base model trained by Lp-Reg from math reasoning. (Part 3)